Circular No. 6235

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET P/1995 S1 = P/1846 D1 (DE VICO)

The following additional precise positions are available:

1995 UT α_{2000} δ_{2000} m_{1} Observer
Sept.20.80420 9^{h}14^{m}32.^{s}68 +0^{o}28′45′′.6  Yusa
20.80660 9 14 33.30 +0 28 53.1 6.5 Otomo
20.80709 9 14 33.36 +0 28 53.7  Yusa
20.80747 9 14 33.58 +0 28 55.7  Otomo
20.82017 9 14 36.38 +0 29 39.6  Yusa
21.74909 9 18 17.40 +1 26 50.2  Garradd
21.75166 9 18 17.95 +1 26 59.6  _{”}
21.75371 9 18 18.48 +1 27 07.3  _{”}
21.77839 9 18 24.34 +1 28 32.5  Cross
21.79288 9 18 27.82 +1 29 33.7  Garradd
21.79390 9 18 28.08 +1 29 38.5  _{”}
21.79731 9 18 28.92 +1 29 43.3  Cross
21.81362 9 18 32.87 +1 30 43.9  _{”}
22.15762 9 19 57.10 +1 52 11.0  Tichý
22.15837 9 19 57.20 +1 52 13.4  _{”}
22.15910 9 19 57.46 +1 52 16.2  _{”}

T. Yusa (Kogota). Communicated by A. Nakamura, Kuma, Ehime, Japan. S. Otomo (Kiyosato). 0.25-m f/3.4 reflector. Comm. by A. Nakamura. G. J. Garradd (Loomberah, N.S.W.). 0.25-m f/4.1 reflector + CCD. K. Cross (Sendai Astronomical Observatory). Comm. by A. Nakamura. M. Tichý, Z. Moravec (Kleť). 0.57-m refl. + CCD + R filter. m_{2} = 11.3.

The following orbital elements by the undersigned are from 61 observations, 1846 Mar. 4-1995 Sept. 22 (mean residual 2′′.2; cf. IAUC 6232):

Epoch = 1995 Oct. 10.0 TT

T = 1995 Oct. 6.0235 TT  ω = 12.^{o}9708
e = 0.962734  Ω = 79.6210 }2000.0
q = 0.658959 AU  i = 85.3981

a = 17.682741 AU n^{o} = 0.0132550 P = 74.357 years The elements for epoch 1846 Mar. 5.0 TT are T = 1846 Mar. 6.044 TT, e = 0.96302, q = 0.66384 AU, ω = 12^{o}.907, Ω = 79^{o}.703, i = 85^{o}.118; these also suggest T = 1922 Apr. 8.4 TT. The ephemerides on IAUC 6232–6233 agree with those computed from this orbit to within ∼ 0′.5 on Oct. 30.

1995 September 22 Daniel W. E. Green